OpenStudy (anonymous):

Select one of the factors of x^3 + 4x2 + 9x + 36 a. none of the above b.(x^2 + 4) c.(x^2 + 9) d.(x + 9)

OpenStudy (anonymous):

This factors to : (x+4)(x^2+9), so C

OpenStudy (anonymous):

how did you do that? can you please explain

OpenStudy (anonymous):

yeah no problem

OpenStudy (anonymous):

factor the first two terms x^3 + 4x2 = x^2(x + 4) Can you factor the second two terms? 9x + 36 = ???

OpenStudy (anonymous):

yeah is 4

OpenStudy (anonymous):

we can factor by grouping: x^3 + 4x2 + 9x + 36 (x^3+4x^2)+(9x+36) x^2(x+4)+9(x+4) (x^2+9)(x+4)

OpenStudy (anonymous):

4? or 9(x + 4)

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

how did the 9 come before the 4? (x^2+9)(x+4)

OpenStudy (anonymous):

well since we grouped this up into: (x^3+4x^2)+(9x+36) we had to factor each one of there terms: For (x^3+4x^2) we factored a x^2 giving us: x^2(x+4) For (9x+36) we factored a 9 giving us: 9(x+4)

OpenStudy (anonymous):

ok thank you ! :)